OpenStudy (anonymous):

I do not understand when to use eitherSin, Cos, Tan or their transverse on a problem.

OpenStudy (anonymous):

Consider the definitions of the different functions and figure out which one fits the situation you have.

OpenStudy (amistre64):

thats what it amounts to, use the one the is best defined in the situaiton

OpenStudy (amistre64):

 you tend to be given a problem that asks you to find one of these values

OpenStudy (amistre64):

sin = y/r cos = x/r tan = y/x

OpenStudy (anonymous):

I have the definitions of them, but a problem like this gives answers with all of them and I dont know whcih on it is

OpenStudy (amistre64):

thats just a picture, whats the question that goes with it?

OpenStudy (anonymous):

What is the length of the piece of wood required to make the side AB of the roof truss?

OpenStudy (amistre64):

so we are given "r" and "t" and asked to find a missing side; its either going to be cos or sin

OpenStudy (amistre64):

since AB is next to, or adjcent, to the angle, that sounds like a cosine to me

OpenStudy (amistre64):

cos(45) = (AB)/12

OpenStudy (amistre64):

there is of course another way to view the problem, but it doesnt give you the practice for your sincostans

OpenStudy (amistre64):